OpenStudy (anonymous):

solve each proportion. 3 over 7 equals 12 over x

OpenStudy (mathmale):

\[3 \over 7 equals 12 \over x ~becomes\] \[\frac{ 3 }{ 7 }=\frac{ 12 }{ x }\]

OpenStudy (mathmale):

How would you solve this equation for x? I'd suggest that you cross multiply as a first step. Try this, please.

OpenStudy (anonymous):

okay I got \[\frac{ 3x }{ 84 }\] Is this correct? If so what is my next step?

OpenStudy (mathmale):

Cross multiplication should have given you the new equation 3x=84, not the ratio 3x/84. Solve 3x=84 for x by dividing both sides by 3: 3x 84 -- = ______ = ? 3 3